• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Chang

    Log entry time 16:05:12 on April19,2001

    Entry number 60826

    keyword=Run Summary

    Run# Start Stop #Events Comments
    3686 Harp run #1263, x=0, y=0
    3687 Harp run #1264, x=0, y=+2
    3689 Harp run #1265, x=0, y=-2
    3690 Harp run #1266, x=-2, y=-2
    3691 Harp run #1267, x=-2, y=0
    3693 Harp run #1268, x=-2, y=+2
    3694 Harp run #1269, x=+2, y=+2
    3695 Harp run #1270, x=+2, y=0
    3696 Harp run #1271, x=+2, y=-2
    3697 14:05 74320 9-foil 12C, forgot to set beam at x=+1
    3698 14:25 14:28 30624 15cm LH2, forgot to set beam at x=+1
    3699 14:32 14:39 125311 Repeat of run #3698 with x=+1, forgot to
    download new trigger file.
    3700 14:47 14:52 108406 Repeat of run #3697 with x=+1, forgot to
    download new trigger file.


    09:05 Still ramping down right arm spectrometer current. As soon as the current
    is down to zero we will then the polarity of both spectrometers. While we
    are waiting for the current to ramp down, Doug is doing arc measurement.
    10:30 Right arm dipole current is zeroed now. Doug called RECON and Scot
    Spiegel. They are on the way to make hall access and change the polarity
    of both spectrometers.
    11:40 Done with the polarity changes. Set both spectrometers to 2.3770 GeV/c.
    11:53 Call MCC to resume beam.
    12:10 Target motion hanged. Reboot target IOC.
    12:40 Doing bulleye harp scans with raster off, 5 uA.
    13:30 Done with bulleye harp scan runs. Right arm dipole is not quite ready yet.
    13:46 MCC informed us that they have injector problem. It will take about 15
    minutes. Right arm dipole is now ready.
    14:00 Beam is back.
    14::55 Change to new kinematics.